OpenStudy (anonymous):

Sugar contains 15 calories per teaspoon. How many calories are in 3/4 cup of sugar?

OpenStudy (decarr432):

How many teaspoons in a cup?

563blackghost (563blackghost):

First determine how many teaspoons are in one cup...

563blackghost (563blackghost):

also

563blackghost (563blackghost):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (decarr432):

Always beat me lol

563blackghost (563blackghost):

^^ its cause I have already ready in a saved notepad for openstudy welcomes...xD

563blackghost (563blackghost):

@Victoria05 are you here? If so please cooperate so @Decarr432 and I can help you with your question ^^

OpenStudy (anonymous):

There are 48 teaspoons in 1 cup

563blackghost (563blackghost):

Correct ^^ so we know of how many teaspoons in one cup now we must multiply 48 by 15 to get how many calories in one cup...\[48 tps \times 15 calories = total||calories\]

OpenStudy (anonymous):

720

563blackghost (563blackghost):

Correct ^^ so now we would multiply this number to 3/4 to get how much calories there are in 3/4 of a cup...so\[720 \div \frac{ 3 }{ 4 }\] or you can change 3/4 to .75 and still multiply the same way...

OpenStudy (anonymous):

60?

563blackghost (563blackghost):

Not quite. First we would multply 720 and 3...\[720\times3=\frac{ total }{ 4 }\]

OpenStudy (anonymous):

Oh okay! 540?

563blackghost (563blackghost):

Correct ^^ So there are 540 calories in 3/4 cup xD Great Job!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

3/4 a cup = 36 teaspoons which I hope helps you since now you can learn the answer which is 540 since 3/4 a cup is 36 teaspoons and a teaspoon is 15 calories. So 15 x 36 = 540. I hope this helps you

563blackghost (563blackghost):

Your welcome ^^

OpenStudy (anonymous):

medal?

OpenStudy (decarr432):

Black should get it